OpenStudy (anonymous):

13. Choose the ordered pair that is the solution to the system of equations. -5y + 4x = 8 -15y + 3x = -3 (3,4/5) (-3,-4) (-3,4/3) (3,0)

OpenStudy (angela210793):

-5y + 4x = 8 (e shumezojme me 3)----->-15y+12x=24 (i zbresim ane per ane) -15y + 3x = -3 -15y+3x=-3 9x=27 x=27/9=3 y---> -5y+12=8 -5y=-4 y=4/5 zgjidhja eshte (3;4/5)

OpenStudy (anonymous):

15y-12x=-24 -15y+3x=-3 =>-9x=-27 => x=3 y= 4/5

OpenStudy (angela210793):

E kape?

OpenStudy (anonymous):

hmm, e kuptova menyren falemnderit shume

OpenStudy (angela210793):

yw :)

OpenStudy (anonymous):

15. Choose the ordered pair that represents the solution to the system of equations. x + y = 4 y = 2x + 1 (1,3) (-1,3) (1,-3) (-1,-3)

OpenStudy (anonymous):

x+y=4 -y+2x=-1 =>3x=3 => x=1 y=3

OpenStudy (angela210793):

(1;3) x + y = 4 --->x+y=4(shumezojme me 2)---->2x+2y=8 (mbeldhim ane per ane) y = 2x + 1 y-2x=1 y-2x=1 3y=9 y=3 x+3=4 x=1

OpenStudy (anonymous):

system: -5y + 4x = 8 -15y + 3x = -3 equal: (3, 4/5)